Ray and Heather are upgrading the storage shed in their backyard. The existing shed has


the shape of a rectangular prism with dimensions 1 m by 2 m by 4 m. By what amount


should they increase each dimension so that the new shed is 9 times the volume of the


original shed? Provide a full algebraic solution.


1
Expert's answer
2021-11-24T06:39:52-0500

Define the following for the dimensions of the rectangular prism.

"w=width, \space l=length,\space h=height"

The volume of the original prism is given as,

"V_1=l_1\times w_1 \times h_1=1\times2\times 4=8m^3"

The volume of the new rectangular prism should be 9 times the volume of the original rectangular prism. Therefore, volume of the new rectangular prism is,

"V_2=V_1\times 9=8\times9=72."

Let "a" be the increase made for each dimension. The dimensions for the new rectangular prism are given as,

"l_2=l_1+a=1+a"

"w_2=w_1+a=2+a"

"h_2=h_1+a=4+a"

The volume of the new rectangular prism is,

"V_2=l_2\times w_2\times h_2=(1+a)\times(2+a)\times(4+a)=72......(i)"

Opening the brackets for equation "(i)" above gives,

"a^3+7a^2+14a+8=72...........(ii)"

We shall solve for the value of "a" in equation "(ii)" as described below.

Let "a=0" then equation "(ii)" gives,

"(0)^3+7(0)^2+14(0)+8=8\not=72". Thus , when "a=0" equation "(ii)" does not hold.

Let "a=1" then equation "(ii)" gives,

"(1)^3+7(1)^2+14(1)+8=30\not=72". Thus , when "a=1" equation "(ii)" does not hold.

Let "a=2" then equation "(ii)" gives,

"(2)^3+7(2)^2+14(2)+8=8+28+28+8=72". Thus, when "a=2" equation "(ii)" holds.

Therefore, each dimension should be increased by "a=2\space m" so that the new shed is 9 times the volume of the original shed.